# Context-Enhanced LLM-Based Framework for Automatic Test Refactoring

### v1

@@ 4.3. RQ2: How effective is the refactoring of UTRefactor?
We are next interested in evaluating the extent to which UTRefactor eliminates bad smells in refactored test code that still passes execution, thus validating the effectiveness of UTRefactor in test refactoring. Additionally, we compare and analyze UTRefactor’s effectiveness against two baseline tools. To assess refactoring effectiveness, we focus on two key metrics in this research question (RQ): the number of passing tests after refactoring and the number of remaining smells in those tests. We also measure the change in the total number of smells before and after refactoring, the greater the reduction rate, the more effective the automatic refactoring is in reducing overall test smells in the project. Table 4 shows the effectiveness of UTRefactor and the two baseline tools in refactoring tests across six Java open-source projects. Out of the 2,375 test smells present across the six projects, UTRefactor reduced the number to 265 after refactoring, achieving an average reduction rate of 89%. The Commons-compress project shows the highest reduction rate, reaching 94%. This demonstrates that UTRefactor is highly effective in reducing test smells. In particular, for the most prevalent smell, AR (Assertion Roulette), UTRefactor had the most significant impact, completely eliminating this smell by successfully adding descriptive messages to all assertions that lacked them. The effectiveness of UTRefactor in addressing different types of test smells will be discussed further in RQ3.
@@
Comparison with Llama3-70B. It is evident that when only providing the test code and corresponding smell types to Llama3-70B for refactoring, it reduces a total of 1,080 smells across the six projects, with an average reduction rate of 55%. This is lower than that achieved by UTRefactor. This shows that while the LLM has the ability to refactor test smells, its effectiveness is limited. After manual inspection of the LLM-refactored tests, we find that it performs best in handling AR (Assertion Roulette) smells. In tests with AR smells, the LLM successfully adds descriptive messages to assertions that lack them, which is attributed to its inherent code understanding capabilities. However, it does not eliminate all AR smells. When multiple smells are present in a test, using the default prompt (as shown in Figure 6) often leads to incomplete smell elimination, leaving AR smells unresolved. This issue is mitigated in UTRefactor by using a checkpoint mechanism, which ensures that each smell is checked and eliminated after refactoring. For other types of smells, the default LLM’s ability is also limited. For example, in handling exception-related smells, without the structured refactoring DSL rules designed in UTRefactor, the Llama3-70B often tends to simply replacing throws statements with try-catch blocks, which does not effectively remove the smell. Overall, the default LLM exhibits more randomness in test refactoring. This characteristic becomes more prominent when multiple smells are present in a single test, resulting in unpredictable refactoring outcomes. In contrast, UTRefactor, with its integration of external knowledge and structured refactoring rules, consistently delivers more effective and stable smell elimination.
@@ 4.4. RQ3: How effective is UTRefactor in eliminating each category of Test Smell?
In this research question, we focus on UTRefactor’s ability to eliminate different types of test smells. To do this, we first detect and collect all types of test smells present in the six open-source projects. We then analyze how effectively UTRefactor removes these smells. As noted in RQ2, UTRefactor successfully reduces the total number of test smells from 2,375 to 265. Figure 8 provides detailed data on UTRefactor’s ability to eliminate each type of test smell. From the figure, we observe that UTRefactor completely eliminates the most prevalent type, i.e., AR (Assertion Roulette) smell. This success is due to UTRefactor’s integration of LLM code understanding capabilities, enabling it to infer appropriate descriptive messages based on the specific assert statements and surrounding code context. These messages help developers quickly identify the cause of errors when tests fail. Additionally, UTRefactor shows significant refactoring results for the two other high-frequency smells, ECT (Exception Catching Throwing) and ETa (Eager Test), reducing their numbers from 526 and 509 to 54 and 20, respectively.
@@
For the cases where smell elimination failed, we conduct a manual review and analysis. From the Figure 8, we observe that UTRefactor’s ability to handle certain less frequent smells is somewhat limited. A notable example is SE (Sensitive Equality), which occurs when the default toString method is used in assertions. The refactoring approach for this smell requires either the project to have overridden the toString method or to provide an equivalent method for object comparison. In other words, if the project neither overrides toString nor offers a functionally equivalent method, UTRefactor is unable to remove this type of smell. Overall, across the six open-source projects, UTRefactor shows less effectiveness in addressing this smell type, primarily due to the lack of overridden toString methods or alternative methods in the project.
@@ 4.5. RQ4: What is the time efficiency of UTRefactor in test refactoring?
In this research question, we focus on the time efficiency of UTRefactor during the test refactoring process. We first collect the time data for each step of UTRefactor and then compare the total time with the two baseline tools. Table 5 illustrates the time spent by UTRefactor at each refactoring step. Across all six open-source projects, the total time from detecting test smells to refactoring the 879 tests with identified smells is 4,379 seconds, averaging 3.8 seconds per test. This highlights UTRefactor’s efficiency in terms of time consumption. Moreover, to accommodate various real-world refactoring scenarios, such as automatically refactoring test smells immediately after developers write a test or refactoring the entire test code of an existing project, UTRefactor supports refactoring at different levels: a single test, a single test file, or all test files within a project.
@@
Next, we analyze the time consumption of UTRefactor. As shown in the Table 5, the most time-consuming step is refactoring the test code—accounting for 87.2% of the total time. This is because UTRefactor relies on LLM API calls to refactor the code, which introduces network latency. Additionally, after receiving the LLM output, we further process the results to facilitate developer evaluation and use of the refactored tests. This includes extracting the refactored test code and merging it into a standalone, executable test file within the target project. The second step is the shortest, as it involves constructing external knowledge, which is a one-time process. Once this external information is built, it can be quickly accessed and loaded during the refactoring process, significantly reducing time consumption at this step.
Figure 9. Comparison of time consumption between UTRefactor and two baseline tools across each project.
Additionally, we compare the time consumption of UTRefactor with the two baselines, as shown in Figure 9. The most time-efficient tool is TESTAXE, which takes 600 seconds, followed by Llama3-70B with a total of 3,291 seconds. TESTAXE is rule-based, and its refactoring process involves parsing the source code into an abstract syntax tree (AST), performing nine types of operations such as node insertion, replacement, and deletion during tree traversal, and finally converting the updated AST back into test code. Since this process does not involve network API calls and the scope of its refactoring rules is limited, it has a relative advantage in terms of time efficiency. Llama3-70B, similar to UTRefactor, requires network calls during the refactoring process. However, Llama3-70B does not involve preprocessing the tests or loading external knowledge, which results in slightly lower time consumption compared to UTRefactor.
@@ 5.1. Strengths of UTRefactor
In UTRefactor, we explore how LLMs perform automatic test refactoring tasks, guided by rules described in DSLs. This approach simulates the process that developers or testers follow during refactoring, from understanding the test’s intent and identifying test smells to following step-by-step instructions outlined in the DSL. Compared to directly using an LLM for test refactoring, UTRefactor significantly reduces the risk of random errors caused by hallucinations, which can make the refactoring process unpredictable. Moreover, if developers directly use an LLM to refactor tests, in addition to the limitations mentioned earlier, there is often incomplete test smell elimination, requiring further manual intervention. In contrast, UTRefactor’s checkpoint mechanism ensures a more thorough refactoring process, especially when multiple test smells are present, greatly reducing the need for manual involvement and lowering associated costs.
@@ 5.2. Integration of UTRefactor with LLMs
We manually review the test refactoring results using GPT-4o. Without guidance from external knowledge on test smells, the default GPT-4o encounters issues in eliminating smells. For example, in the testFindRangeBounds of the jfreechart, where four types of smells are present. GPT-4o removes only three smells but fails to eliminate the most common smell, Assertion Roulette. In other cases, we observe that GPT-4o attempts to address Assertion Roulette by adding comments to assertion statements instead of explicitly adding a message, which does not properly resolve the smell. In contrast, UTRefactor leverages its DSL to clearly instruct the addition of a message to assertions and ensure that all smells are addressed during refactoring, avoiding the issues caused by the default LLM approach. In another example, the testDrawWithNullInfo contains a complex try/catch structure detected as Exception Catching Throwing. GPT-4o simply removes this block and adds an explicit throw in the method signature, but this fails to eliminate the smell. By contrast, guided by UTRefactor’s DSL, the tool uses assertNotThrow to replace the try/catch block and successfully resolves the smell. This demonstrates that UTRefactor effectively eliminates smells even when using other LLMs like GPT-4o.

### v2

@@ 4.3. RQ2: How effective is the refactoring of UTRefactor?
We are next interested in evaluating the extent to which UTRefactor eliminates bad smells in refactored test code that still passes execution, thus validating the effectiveness of UTRefactor in test refactoring. Additionally, we compare and analyze UTRefactor’s effectiveness against two baseline tools. To assess refactoring effectiveness, we focus on two key metrics in this research question (RQ): the number of passing tests after refactoring and the number of remaining smells in those tests. We also measure the change in the total number of smells before and after refactoring. A higher reduction rate indicates greater effectiveness of our automated refactoring approach in minimizing test smells across the project.
Table 5 shows the effectiveness of UTRefactor and the two baseline tools in refactoring tests across six Java open-source projects. Out of the 2,375 test smells present across the six projects, UTRefactor reduced the number to 265 after refactoring, achieving an average reduction rate of 89%. The Commons-compress project shows the highest reduction rate, reaching 94%. This demonstrates that UTRefactor is highly effective in reducing test smells. In particular, for the most prevalent smell, AR (Assertion Roulette), UTRefactor had the most significant impact, completely eliminating this smell by successfully adding descriptive messages to all assertions that lacked them. The effectiveness of UTRefactor in addressing different types of test smells will be discussed further in RQ3.
@@
Comparison with Llama3-70B. It is evident that when only providing the test code and corresponding smell types to Llama3-70B for refactoring, it reduces a total of 1,080 smells across the six projects, with an average reduction rate of 55%. This is lower than that achieved by UTRefactor. This shows that while the LLM has the ability to refactor test smells, its effectiveness is limited. After manual inspection of the LLM-refactored tests, we find that it performs best in handling AR (Assertion Roulette) smells. In tests with AR smells, the LLM successfully adds descriptive messages to assertions that lack them, which is attributed to its inherent code understanding capabilities. However, it does not eliminate all AR smells. When multiple smells are present in a test, using the default prompt (as shown in Figure 6) often leads to incomplete smell elimination, leaving AR smells unresolved. This issue is mitigated in UTRefactor by using a checkpoint mechanism, which ensures that each smell is checked and eliminated after refactoring. For other types of smells, the default LLM’s ability is also limited. For example, in handling exception-related smells, without the structured refactoring DSL rules designed in UTRefactor, the Llama3-70B often tends to simply replace throws statements with try-catch blocks, which does not effectively remove the smell. Overall, the default LLM exhibits more randomness in test refactoring. This characteristic becomes more prominent when multiple smells are present in a single test, resulting in unpredictable refactoring outcomes. In contrast, UTRefactor, with its integration of external knowledge and structured refactoring rules, consistently delivers more effective and stable smell elimination.
Effectiveness of Assertion Roulette Elimination. Adding any string message to assertions can eliminate the Assertion Roulette (AR) smell. However, the generated messages should clearly convey the intent of the assertion. To evaluate the quality of these messages, we conduct an experiment using GPT-4o, which is tasked with evaluating messages based on two criteria: ❶ Is the message clear and explicit?Does the message align with the intent of the assertion? In this experiment, we collect all 1,230 generated assertion messages. The results show that only 43 messages (3.5%) are flagged as Unclear, and all of these messages fail to meet criterion ❶ (clarity of the message). After manual review, we find that these cases share a common feature: they lack sufficient context. For example, numerical variables (such as Magic Numbers) are extracted as constants with meaningful names. assertEquals(expectedDatasetIndex, li.getDatasetIndex(), "The dataset index of the legend item should be 1"), GPT-4o marks the message as unclear because it misinterprets expectedDatasetIndex as a variable rather than a constant. However, the generated message is reasonable, as the intent behind it is clear. Therefore, we conclude that LLMs can generate clear and descriptive assertion messages that align with the intent of the assertion.
Ablation of the DSL. We conduct an ablation study by comparing the effectiveness of UTRefactor with a version excluding the DSL, while keeping all other aspects the same (e.g., detailed descriptions of test smells, test context, and the smell checkpoint mechanism). The results show that without the DSL the reduction rate of test smells is 67%, which is an improvement over using LLM alone (55%) but significantly lower than the 89% achieved with the DSL.
Additionally, the natural language descriptions of test smells allow the LLM to partially address them effectively. For example, in 509 cases of Exception Catching Throwing, 175 cases (34%) successfully eliminate the smell using assertDoesNotThrow or assertThrow. However, the remaining cases show randomness, such as removing try-catch blocks or replacing throws with try-catch. In contrast, the DSL clearly defines rules for eliminating such smells, covering different scenarios (e.g., adding exception descriptions), leading to more consistent results.
Finally, we analyze all refactored tests that fail at runtime after successful compilation. Most failures occur when handling Duplicate Assertions, as our DSL refactors the original tests into JUnit 5 parameterized tests. The LLM occasionally introduces errors in the parameterized data provided via @CsvSource, causing runtime failures. We find that the LLM struggles with converting tests into the parameterized format.
@@ 4.4. RQ3: How effective is UTRefactor in eliminating each category of Test Smell?
In this research question, we focus on UTRefactor’s ability to eliminate different types of test smells. Refactoring based on the type of code smells can lead to the same test being refactored multiple times, as many smells often coexist within a single test. Additionally, the order in which refactoring is performed can influence the final outcome (sec 3.2.3). To address these challenges, our refactoring process targets individual tests and simultaneously eliminates all smells. We first detect and collect all types of test smells present in the six open-source projects. We then analyze how effectively UTRefactor removes these smells.
As noted in RQ2, UTRefactor successfully reduces the total number of test smells from 2,375 to 265. Figure 8 provides detailed data on UTRefactor’s ability to eliminate each type of test smell. From the figure, we observe that UTRefactor completely eliminates the most prevalent type, i.e., AR (Assertion Roulette) smell. This success is due to UTRefactor’s integration of LLM code understanding capabilities, enabling it to infer appropriate descriptive messages based on the specific assert statements and surrounding code context. These messages help developers quickly identify the cause of errors when tests fail. Additionally, UTRefactor shows significant refactoring results for the two other high-frequency smells, ECT (Exception Catching Throwing) and ETa (Eager Test), reducing their numbers from 526 and 509 to 54 and 20, respectively.
@@
Figure 9. Comparison of time costs between UTRefactor and two baseline tools across each project.
For the cases where smell elimination failed, we conducted a manual review and analysis. From the Figure 8, UTRefactor’s ability to handle certain less frequent smells is somewhat limited. A notable example is SE (Sensitive Equality), which occurs when the default toString method is used in assertions. The refactoring approach for this smell requires either the project to have overridden the toString method or to provide an equivalent method for object comparison. In other words, if the project neither overrides toString nor offers a functionally equivalent method, UTRefactor is unable to remove this type of smell. Overall, across the six open-source projects, UTRefactor shows less effectiveness in addressing this smell type, primarily due to the lack of overridden toString methods or alternative methods in the project.
@@ 4.5. RQ4: What is the time efficiency of UTRefactor in test refactoring?
In this research question, we focus on the time efficiency of UTRefactor during the test refactoring process. We first collect the time data for each step of UTRefactor and then compare the total time with the two baseline tools.
@@
Table 6 illustrates the time spent by UTRefactor at each refactoring step. Across all six open-source projects, the total time from detecting test smells to refactoring the 879 tests with identified smells is 4,379 seconds, averaging 3.8 seconds per test. This highlights UTRefactor’s efficiency in terms of time consumption. Moreover, to accommodate various real-world refactoring scenarios, such as automatically refactoring test smells immediately after developers write a test or refactoring the entire test code of an existing project, UTRefactor supports refactoring at different levels: a single test, a single test file, or all test files within a project.
@@ 5.1. Strengths of UTRefactor
In UTRefactor, we explore how LLMs perform automatic test refactoring tasks, guided by rules described in DSLs. This approach simulates the process that developers or testers follow during refactoring, from understanding the test’s intent and identifying test smells to following step-by-step instructions outlined in the DSL. Compared to directly using an LLM for test refactoring, UTRefactor significantly reduces the risk of random errors, which can make the refactoring process unpredictable.
To further illustrate this issue, we randomly select 100 cases where the LLM fails to eliminate smells and conduct an analysis. Among them, 65 cases fail due to the LLM’s arbitrary refactoring behavior, such as removing throws declarations or try-catch blocks without resolving the Exception Catching Throwing smell. 22 cases fail due to hallucinations during the elimination of the Eager Test smell, where the LLM introduces non-existent functions or variables into split tests. 23 cases fail because the LLM generates helper methods with empty bodies and comments instructing users to complete them.
@@ 5.2. Integration of UTRefactor with LLMs
We manually review the test refactoring results using GPT-4o. Without guidance from external knowledge on test smells, the default GPT-4o encounters issues in eliminating smells. For example, in the testFindRangeBounds of the jfreechart, where four types of smells are present. GPT-4o removes only three smells but fails to eliminate the most common smell, Assertion Roulette. In other cases, we observe that GPT-4o attempts to address Assertion Roulette by adding comments to assertion statements instead of explicitly adding a message, which does not properly resolve the smell.
In contrast, UTRefactor leverages its DSL to clearly instruct the addition of a message to assertions and ensure that all smells are addressed during refactoring, avoiding the issues caused by the default LLM approach. In another example, the testDrawWithNullInfo contains a complex try/catch structure detected as Exception Catching Throwing. GPT-4o simply removes this block and adds an explicit throw in the method signature, but this fails to eliminate the smell. By contrast, guided by UTRefactor’s DSL, the tool uses assertNotThrow to replace the try/catch block and successfully resolves the smell. This demonstrates that UTRefactor effectively eliminates smells even when using other LLMs like GPT-4o.
@@ 7. Conclusion and Future Work
8. Data Availability
Our tool is available at (utr, [n. d.]).
Acknowledgements.
This work was supported by the National Key R&D Program of China (No. 2024YFB4506400) and sponsored by CCF-Huawei Populus Grove Fund.